# Communication and Energy-Aware Multi-UAV Coverage Path Planning for Networked Operations

### v1

@@ 7                   $R_{req}^{{min}}\leftarrow R_{req};\ \{\Delta s_{i}\}_{best}\leftarrow\{\Delta s_{i}\}$;
9
10
11
Result: Return $\{\Delta s_{i}\}_{best}$ and $R_{req}^{{min}}$
Algorithm 2 Optimize launch points for connectivity
III-D Estimating Energy Consumption for Paths
Estimating the energy consumption of UAV flight paths can be performed using various methods, depending on the desired accuracy level. For instance, in [26], the authors present a method that calculates UAV energy consumption based on comprehensive UAV parameterization. Alternatively, [toppra] describes an approach that estimates energy consumption by generating trajectories of the flight path.
While these methods provide accurate energy consumption estimates, they often require detailed UAV parameters or complex trajectory generation processes. In our work, we prioritize a more lightweight and computationally efficient approach that can be easily integrated into the optimization process without compromising the overall accuracy of the energy consumption estimates.
To achieve this, we propose a method for estimating energy consumption based on analyzing the path waypoints, power consumption model, and speed parameters of the mission. This approach allows us to capture the essential factors influencing energy consumption while maintaining a simple and efficient calculation process. our power consumption model is defined as follows:
$P(v)=P_{\text{b}}+P_{\text{d}}(s)$ (5)
Where:
  • $P_{\text{b}}$: power required to run electronics, sensors, and communication systems.
  • $P_{\text{d}}(s)$: power required to overcome air resistance while moving with speed $s$, calculated separately for straight paths and turns
Path segmentation and energy estimation: Each UAV path is divided into straight-line and turning segments by considering the turning radius at each corner waypoint. This segmentation helps identify the total time spent on straight-line movements ($\Delta T_{s_{f}}$) and turning maneuvers ($\Delta T_{s_{t}}$). For each path, we calculate the energy consumption using the UAV’s empirically derived power model, which distinguishes between forward flight ($P_{d}(s_{f})$) and turning ($P_{d}(s_{t})$) power requirements. The total energy consumption for a path is then calculated as:
$E_{\text{path}}=P_{\text{b}}\cdot T+P_{d}(s_{f})\cdot\Delta T_{s_{f}}+P_{d}(s_{t})\cdot\Delta T_{s_{t}}$ (6)
DARP paths consistently produce rectangular turns, which allows us to assume uniform energy consumption for all turns. This characteristic simplifies the energy estimation process, enabling our algorithm to rapidly calculate the energy requirements for each UAV path using the power consumption model. The total energy requirement for the entire mission is then determined by summing the energy consumption of each individual path. This streamlined approach, enables the efficient evaluation of numerous potential paths while maintaining adequate accuracy for comparative assessment.
@@ III-D Estimating Energy Consumption for Paths
III-E Optimization Problem
The optimization problem involves finding the optimal DARP initialization positions for each UAV on a 2D grid that minimize both the required communication range and energy consumption. Given a grid with $C$ cells and $N$ UAVs, the search space size is $C^{N}$, growing exponentially and making exhaustive search computationally infeasible. Therefore, we employ Bayesian optimization, specifically the Tree-Structured Parzen Estimator (TPE) algorithm implemented within the Optuna framework, to efficiently explore this vast search space. TPE builds probability models of the objective function values conditioned on hyperparameter configurations and uses them to determine the most promising areas to sample. The objective function $f_{o}$ is defined as:
$f_{o}=R_{req}^{{min}}+\lambda\cdot E_{tot}$ (7)
where $R_{req}^{{min}}$ is the minimum required communication range throughout the mission (in meters), $E_{tot}$ is the total estimated energy consumption (in watt-hours), and $\lambda$ is the user-defined trade-off coefficient. The optimization is subject to constraints of complete area coverage and avoidance of NFZs. Optuna’s implementation allows for efficient parallel computing and dynamic pruning of unpromising trials, significantly speeding up the optimization process. Here, an ‘unpromising trial’ is defined as one whose objective function value is significantly worse than the current best, allowing early termination to save computational resources.
Decision variables: Key variables influence the optimization process:
  • Startup Trials: Initial random trials build a model of the search space, providing a baseline for Bayesian optimization. We use 20% of the total trials as startup trials to establish this baseline.
  • Pruning Mechanism: A median pruner is used to terminate unpromising trials early. If a trial’s intermediate objective value is worse than the median of previous trials’ values at the same step, it is pruned.
  • $N_{\text{ei\_candidates}}$: Controls the number of candidates sampled in each TPE iteration, balancing thorough exploration with computation time. This choice impacts the trade-off between exploration and exploitation.
III-F Special Case: Formation Flying
Formation flying becomes essential when the inter-UAV communication range is significantly limited (3-5 times the path width). However, the rigid nature of formation shapes makes it challenging to implement in complex ROIs, particularly those with NFZs. A basic model of the formation flying strategy for large-area coverage is discussed in [25], but it lacks consideration for formation shape changes during turns and is not optimized for energy efficiency. For large areas within regular polygonal ROIs and scenarios with strict communication constraints, we propose a communication-aware formation path generation method, as outlined in Algorithm 3. In this approach, the combined formation footprint $F_{c}$ is calculated by summing the widths of individual UAV footprints and treated as a single footprint for generating STC path. Offset paths are then created for individual UAVs with parallel spacing corresponding to their respective footprints, followed by a connectivity assessment performed after time sampling and starting point optimization.
Input: ROI, $N$, Sensor footprints $F_{u_{i}}$
Output: Optimized formation paths $P_{u_{i}}$ for each UAV
1 Discretize ROI into grid cells;
2 Align grid via shifts and rotations;
3 Generate Combined Coverage Path
4 Calculate formation width: $F_{c}=\sum_{i=1}^{N}F_{u_{i}}$;
5 Generate single STC path $P_{c}$ for $F_{c}$;
6 Generate Individual UAV Paths
7 foreach UAV $i$ from 1 to $N$ do
8       Create offset paths $P_{u_{i}}$ parallel to $P_{c}$;
9
10Time-sample all paths using $s_{f}^{u_{i}}$ $s_{t}^{u_{i}}$;
11 Optimize Starting Positions
12 foreach trial in $N_{shift}$ do
13       Adjust start position and compute $R_{req}$;
14       Minimize $R_{req}$ with optimal shifts;
15
16Result: Save optimized paths $P_{u_{i}}$;
Algorithm 3 Formation Path Planning
III-G Operational Constraints and Assumptions
The following assumptions and constraints were applied in the experiment:
  1. 1.
    Uniform UAV Parameters: All UAVs have identical communication range and communicating with a mesh architecture.
  2. 2.
    Calculated Starting Points: UAVs begin their mission from pre-determined starting points only. They all start the mission simultaneously.
  3. 3.
    Exclusion of Takeoff and Landing: The energy for takeoff and landing phases is excluded from the calculations, focusing solely on the mission’s operational phase.
IV Evaluation results
To evaluate the effectiveness of the proposed approach, we first conducted simulations using the ArduPilot Software-in-the-Loop (SITL) simulator. After confirming its performance in the simulated environment, we carried out real-world tests with three UAVs to assess the method’s real-world accuracy and performance. For benchmarking purposes, we compared our approach with three state-of-the-art methods within the SITL environment, as illustrated in Fig. 2. To the best of our knowledge, our method is the first to generate communication-aware coverage paths for arbitrarily shaped ROIs, including support for NFZs, while also allowing users to balance energy consumption and communication requirements in the optimization process. In contrast, existing methods primarily focus on minimizing energy consumption, either directly or through related metrics. To ensure accurate energy estimation, we developed a power consumption model from real-world experiments with our UAVs, which was then applied to the simulations to closely reflect actual operational conditions.
IV-A Hardware Specifications and Flight Metrics
@@
All simulations were conducted on a computer equipped with a 12th Gen Intel(R) Core(TM) i7-12700K CPU, 32GB RAM, and running Ubuntu 20.04.6 LTS OS. For the experiments, we used a custom-built quadcopter platform, featuring a Raspberry Pi4B as the onboard computer and a Ubiquity Bullet M2 [32] with OpenWRT firmware for wireless mesh networking. The UAV operated at a forward flight speed of 5 m/s, a turning speed of 3 m/s, and a mission altitude of 30 m above ground level. Time-stamped waypoint paths were used to synchronize the relative positions of the UAVs throughout the mission. The basic power consumption $P_{\text{b}}$ was calculated as 5W, with additional power requirements of $P_{\text{drag}}(s_{f})$ = 505W $\pm$ 10W for forward speed and $P_{\text{drag}}(s_{t})$ = 498W $\pm$ 8W for turning speed. All optimization operations were performed with 3000 DARP trials ($N_{darp}$) and 1000 shift trials ($N_{shift}$). With the given computer specifications, these optimizations took approximately 60 minutes to complete.
IV-B Validation in Simulation
@@
The performance of the proposed method was evaluated and compared with existing communication and energy-aware approaches in simulations using three UAVs. The comparison included three existing implementations: EA-mCPP [10], the O-DARP approach [12], and a formation flying strategy [25]. The test scenarios applied to each ROI were:
  • Cape: An ROI used in [28], with the same parameters (35 m coverage footprint) and an area of 903,300 m2.
  • Island: A semi-circular shape with a 15 m coverage footprint and an area of 71,000 m2.
  • Rectangle: A rectangular ROI with a 15 m coverage footprint and an area of 77,000 m2.
  • Simple: A simple polygonal shape with a 15 m coverage footprint and an area of 71,000 m2.
  • Complex: A complex shape with a 15 m coverage footprint and an area of 105,000 m2.
  • Island with NFZ: An island shape with an NFZ, a 15 m coverage footprint, and an area of 70,000 m2.
@@
The visual representation of five of the generated paths is shown in Fig. 3. Additionally, the connectivity estimation for complex and island ROIs with NFZs is explained in the supplementary video.
@@
Figure 3: Scenarios used for experimental evaluation of the proposed method. Shade-of-grey polygons represent arbitrary ROI, and blue polygon is the processed ROI.
@@
Fig. 2 presents the simulation results of the proposed solution across cape, island, and rectangular-shaped regions of interest (ROIs), with varying energy-communication trade-off coefficients ($\lambda$). Analyzing the impact of $\lambda$ across different ROIs shows that while the most energy-efficient coverage paths require a larger connectivity range, the most connected paths are not always the most energy-efficient. In Fig. 2, the Pareto front is marked in orange, indicating the optimal values tuned for different levels of energy and communication awareness. The energy and communication requirements of the energy-aware mCPP (EA-mCPP) method [10] are shown with a green rectangle mark, while the requirements of the optimal DARP (O-DARP) algorithm [12] are indicated with a purple triangle mark. The performance of the formation flying strategy [25] is compared in the rectangular ROI in Fig. 2(c).
@@
IV-C Real World Experiments
To validate the feasibility of the proposed method in real-world scenarios and assess the accuracy of the simulation, we conducted two types of experiments, as discussed in Section IV-A. The first experiment focused on analyzing the power consumption model of the UAV at different phases of the mission. Next, the accuracy of the range requirement calculation and energy estimation technique was verified by performing a multi-UAV cooperative coverage mission with three UAVs under the specified flying conditions. The mission was conducted over the simple polygon ROI shown in Fig. 3(d), with a coverage area of 71,000 m2.
The connectivity strength along the path is shown in a heat map in Fig. 4. Blue arrows mark the vehicles’ starting positions, with strong connections in green and weak connections in red. This visualization helps identify potential weak spots in the network, allowing for mitigation of connectivity issues and enabling more effective data exchange planning.
@@ IV-C Real World Experiments
The communication range requirements and mission duration from the real-world experiment were compared with the estimated results, as shown in Fig. 5. The estimated minimum range requirement was 125.3 meters, while the real-world experiment showed a slightly lower range requirement of 124.21 meters. This minor difference between the estimated and actual paths is due to the fact that in real-world situations, UAVs take curved turns at corners instead of sharp rectangular turns. Similarly, the estimated mission duration was 385 seconds, but the actual mission took 389 seconds. This 4-second difference in mission time is attributed to a small error in the turning speed profile of the estimated and actual drones. In the estimation, we did not consider the acceleration or deceleration for switching between forward speed and turning speed. However, actual UAVs have an acceleration model, which introduces a small error of about 150ms during each turn. This error accumulates to 4 seconds over the entire mission.
The real-world experiment demonstrates that the proposed estimation method achieves 99.9% accuracy in determining the minimum range requirement for the mission (Fig. 5). Minor deviations between the estimated and actual range requirements occur during specific time intervals (155-210s and 300-330s) due to the delay in UAV paths resulting from differences in turning speed, which causes a variation in the relative positions of the UAVs compared to the estimated positions.The difference in energy consumption was small, with the estimation predicting 130.44 kW and the actual energy consumption being 136.10 kW, resulting in a 95.8% accuracy. This discrepancy is primarily due to approximation techniques and external environmental influences.
V CONCLUSION
We proposed a novel communication and energy-aware multi-UAV coverage path planning method based on area discretization and Minimum Spanning Tree (MST) generation. To the best of our knowledge, this is the first solution that generates communication-aware paths for inter-UAV cooperative missions requiring continuous communication. By introducing the communication-energy coefficient in the optimization process, our method uniquely allows users to adjust the trade-off between energy awareness and communication awareness during path generation. We compared our approach to state-of-the-art methods through a series of simulation experiments, where it outperformed them in most cases, reducing communication range requirements by 20-60%. Additionally, our solution demonstrated superior energy awareness compared to existing communication-aware methods. The feasibility of the method was also validated in simulations for arbitrary polygonal shapes and polygons with no-fly zones. Furthermore, we verified the effectiveness of the proposed method in real-world scenarios using a set of three UAVs.

### v2

@@ 7       Use STC on each $\mathcal{R}_{i}$ to form paths;
9      for $j=1$ to $N_{\text{launch}}$ do // Inner loop
10             Generate launch points ${\mathbf{k}}$ via inner TPE;
11             Compute $r(\mathcal{T}(\mathbf{p},\mathbf{k}))$ and prune if too high;
12             Update inner TPE with $({\mathbf{k}},r(\mathcal{T}(\mathbf{p},\mathbf{k})))$;
13
14      $\mathbf{k}(\mathbf{p})\leftarrow\arg\min_{\mathbf{k}}\,r(\mathcal{T}(\mathbf{p},\mathbf{k}))$;
15       Update TPE with $(\mathbf{p},f_{o}(\mathbf{p},{\mathbf{k}}(\mathbf{p})))$;
16
[…]
return $\mathcal{T}^{\star}=\mathcal{T}(\mathbf{p}^{\star},{\mathbf{k}}(\mathbf{p}^{\star}))$
Algorithm 1 Comm-Energy Aware mCPP
Remark 1.
For scenarios with severely constrained communication ranges (typically 3–5 times the sensor footprint width), we develop a formation-flying strategy that ensures continuous connectivity. While formation-based coverage has been explored in [25], existing approaches do not account for formation adaptations during turns or explicit communication optimization.
To handle regular ROIs with tight connectivity limits, define a combined formation footprint $\mathcal{F}_{c}=\bigcup_{i=1}^{N}\mathcal{F}({\mathbf{x}}_{i}(t))$. We use it to generate a single reference trajectory $\mathcal{T}_{c}$ via STC, and keep the individual UAV paths as parallel offsets to the main trajectory, so that, ${\mathbf{x}}_{i}(t)={\mathbf{x}}_{c}(t)+d_{i}\,\mathbf{n}(t)$, where $d_{i}=(2i-N-1)w/2$ is the lateral offset, $w$ is the sensor footprint width, and $\mathbf{n}(t)$ is the normal to the reference path. The connectivity radius $r(\mathcal{T}_{k})$ is optimized as described in Section IV-B. Although restricted to regular-shaped ROIs without NFZs, this approach guarantees connectivity via a maintained formation, making it particularly effective for large-scale missions dominated by severe communication constraints.
@@ 15       Update TPE with $(\mathbf{p},f_{o}(\mathbf{p},{\mathbf{k}}(\mathbf{p})))$;
V Evaluation results
The effectiveness of the proposed mCPP method was evaluated through simulations and real-world tests. The evaluation began with performance validation using the ArduPilot Software-in-the-Loop (SITL) simulator, followed by real-world experiments with three UAVs to assess accuracy and practical performance.
V-A Hardware Specifications and Flight Metrics
@@
All simulations were conducted on a computer equipped with a 12th Gen Intel(R) Core(TM) i7-12700K CPU, 32GB RAM, and running Ubuntu 20.04.6 LTS. For the real world experiments, we used a custom-built quadcopter platform, featuring a Raspberry Pi4B as the onboard computer and a Ubiquity Bullet M2 [34] with OpenWRT firmware for wireless mesh networking. We utilized three UAVs operating at an altitude of 45m for both simulation and real-world testing. The UAV speeds and the estimated energy consumption parameters are shown in Table II. All optimization operations were performed with 3000 DARP trials ($N_{\text{darp}}$) and 1000 shift trials ($N_{\text{launch}}$). With the given computer specifications, these optimizations took approximately 60 minutes to complete.
TABLE II: UAV Power Consumption by Flight State
UAV State Speed (m/s) Power consumption (W)*
Mean Max Min
Forward flight 5 488 510 470
Turning flight 3 509 536 461
Hovering 0 492 510 460
*Measurements taken in normal weather conditions.
V-B Validation in Simulation
@@
The performance of the proposed method was evaluated through simulations using three UAVs. We compared our method with three existing approaches: Energy-Aware mCPP (EA-mCPP) [10], the Optimal DARP (O-DARP) approach [12], and a Formation flying strategy [25]. These comparisons were carried out across multiple ROIs, each varying in complexity and size. The test scenarios included the following ROIs, as detailed in Table III:
TABLE III: ROI Specifications for Simulation Validation
ROI Shape Total Sensor $\Delta t$ (s)
Area (m2) Footprint (m2)
Cape [28] 903,300 1225 5
Island 71,000 225 1
Rectangle 77,000 225 1
Simple 71,000 225 1
Complex 105,000 225 2
Island with NFZs 70,000 225 1
@@
The visual representation of five of the generated paths is shown in Fig. 3. The mission altitude used for the Cape shape ROI is 70m above ground lavel and 45 meters for the rest of the ROIs. Additionally, the connectivity estimation for complex and island ROIs is explained in the supplementary video.
@@
Figure 3: Scenarios used for experimental evaluation of the proposed method. Shade-of-grey polygons represent arbitrary ROI, blue polygon is the processed ROI and solid red polygons are NFZs.
@@
Fig. 2 presents the simulation results of the proposed solution across cape, island with NFZs, and rectangular-shaped ROIs, with varying energy-communication trade-off coefficients ($\lambda$). Analyzing the impact of $\lambda$ across different ROIs shows that while the energy-efficient coverage paths have a higher range requirement, the minimum range required paths are not always energy-efficient. In Fig. 2, the Pareto front is marked in orange, indicating the optimal values tuned for different levels of energy and communication awareness. The energy and communication requirements of the EA-mCPP method are shown with a green rectangle mark, while the requirements of the O-DARP algorithm are indicated with a purple triangle mark. The performance of the formation flying strategy is compared in the rectangular ROI in Fig. 2(c).
@@
V-C Real World Experiments
To validate the feasibility of the proposed method in real-world scenarios and assess the accuracy of the simulation, we conducted a multi-UAV cooperative coverage mission with three UAVs. The mission was conducted over the simple polygon ROI mentioned in Table.
The connectivity strength along the actual flight path is shown in Fig. 4. Blue arrows represent the vehicles’ starting positions, with strong connections in green and weak connections in red dots. This visualization helps identify potential connectivity weak spots in the path, allowing for mitigation of connectivity issues, more effective data exchange planning, and enabling adaptive transmission power control for UAVs to conserve energy. To optimize data traffic at these weak spots, adaptive data compression and proactive data caching can be employed, ensuring that critical information is prioritized and transmitted efficiently, minimizing the impact of connectivity issues throughout the flight path.
@@ V-C Real World Experiments
The communication range requirements and mission duration from the real-world experiment were compared with the estimated results, as shown in Fig. 5. The estimated minimum range requirement was 125.3 meters, while the real-world experiment showed a slightly lower range requirement of 124.21 meters. This minor difference between the estimated and actual paths is due to the fact that in real-world situations, UAVs take curved turns at corners instead of sharp rectangular turns. Similarly, the estimated mission duration was 385 seconds, but the actual mission took 389 seconds. This 4-second difference in mission time is attributed to a small error in the turning speed profile of the estimation and actual drones. In the estimation, we did not consider the acceleration or deceleration for switching between forward speed and turning speed. However, actual UAVs have an acceleration model, which introduces a small error of about 150ms during each turn. This error accumulates to 4 seconds over the entire mission.
The real-world experiment demonstrates that the proposed estimation method achieves 99.9% accuracy in determining the minimum range requirement for the mission (Fig. 5). Minor deviations between the estimated and actual range requirements occur during specific time intervals (155-210s and 300-330s) due to the delay in UAV paths resulting from differences in turning speed, which causes a variation in the relative positions of the UAVs compared to the estimated positions. This discrepancy is primarily due to approximation techniques and external environmental influences.
VI CONCLUSION
We introduce a novel communication and energy-aware multi-UAV coverage path planning method that combines DARP+STC path generation with iterative connectivity assessment optimization. To the best of our knowledge, our approach is the first to generate communication-aware coverage paths for multi-UAV cooperative missions requiring continuous inter-UAV communication. The method incorporates a communication-energy trade-off parameter, allowing users to balance energy and communication awareness during path generation. Simulations demonstrate that our approach outperforms state-of-the-art methods, reducing communication range requirements by 20-60% and exhibiting better energy awareness compared to existing communication-aware method. The method’s feasibility is validated for arbitrary polygonal shapes and polygons with no-fly zones. Real-world experiments with three UAVs confirm the effectiveness of the proposed solution.